SEPT Housing

Stephan Böttcher  Jörg Falenski
Reinhold Müller-Mellin  Eckbert Rode

Christian-Albrechts-Universität zu Kiel

Final DRAFT: October 23, 2002, Version: 1.12, 2002/10/22

Abstract

This note describes the mechanical housings of the Solar Electron Proton Telescope (SEPT) for IMPACT on the NASA space probes STEREO, for the purpose of a structural analysis to verify environmental specifications required for launch.

Contents

1 Description of the Housing 4 1.1 Materials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Electronics box . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.3 Sensor housing . . . . . . . . . . . . . . . . . . . . . . . . . . 7

1.3.1 Magnet assembly . . . . . . . . . . . . . . . . . . . . . 7 1.3.2 Sensor Side support . . . . . . . . . . . . . . . . . . . . 8 1.3.3 Collimators . . . . . . . . . . . . . . . . . . . . . . . . 8 1.3.4 Detector assembly . . . . . . . . . . . . . . . . . . . . 9 1.3.5 The sensor doors . . . . . . . . . . . . . . . . . . . . . 10

1.4 Connectors and Miscellaneous Attachments . . . . . . . . . . . 10 1.4.1 Pinpullers . . . . . . . . . . . . . . . . . . . . . . . . . 10 1.4.2 Purge nipple . . . . . . . . . . . . . . . . . . . . . . . . 12 1.4.3 Detector signals . . . . . . . . . . . . . . . . . . . . . . 12 1.4.4 SEP connection . . . . . . . . . . . . . . . . . . . . . . 13 1.4.5 Thermistor . . . . . . . . . . . . . . . . . . . . . . . . 13 1.4.6 Thermal hardware . . . . . . . . . . . . . . . . . . . . 13 1.4.7 Ground Stud . . . . . . . . . . . . . . . . . . . . . . . 15

2 Vibration levels 15 2.1 Instrument Stiffness . . . . . . . . . . . . . . . . . . . . . . . . 15 2.2 Instrument Sine Sweep Vibration Tests . . . . . . . . . . . . . 15 2.3 Random Vibration Levels . . . . . . . . . . . . . . . . . . . . 15

List of Figures

1 SEPT Housing overview. . . . . . . . . . . . . . . . . . . . . . 5 2 The body of the electronics box. . . . . . . . . . . . . . . . . . 6 3 The electronics box with covers. . . . . . . . . . . . . . . . . . 6 4 The electronics box with bottom (left) and top (right) covers. 6 5 Sensor top view. Silicon detectors (cyan), magnets (magenta),

foil (yellow), foil carrier ring (brown), detector housing (green), Delrin (gray). . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

6 Magnet assembly. Magnets and yokes (magenta), frame and

magnet support (dark blue). . . . . . . . . . . . . . . . . . . . 8 7 Sensor side support. The side supports slide from below be-

tween the extensions of the sensor frame, and the inner support wings of the side supports slide into the slits of the frame and its covers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 8 Sensor side supports and collimator top view. . . . . . . . . . 9 9 Collimator assembly. The positions of the screws are indicated

on the front view. The grayed screw positions are for the reverse side. Some positions are shared. . . . . . . . . . . . . . 9

10 Sensor doors side view. Left: doors in closed position, right:

doors in open position. . . . . . . . . . . . . . . . . . . . . . . 11 11 Sensor doors front view, with the doors opened. The pinpuller

rod is drawn in both the idle (dark blue) and activated (pink) position. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

12 Pinpuller mounting top view. Also shown is the connector for

the supply of the pinpuller activation current. . . . . . . . . . 12 13 Pinpuller mounting side view. Only the near half of the sensor

is shown. The sensor doors are drawn both opened and closed. 12 14 Purge nipple and thermistor (green). The purge nipple is the

upper left and the thermistor is the lower right part in green. . 13 15 Cables and connectors front view. The detector signal cables

are drawn in light green. The connector on the right between the support wings of the E-box supplies the thermal hardware. The MDM connector on the front is the link to SEP central. To its left is the grounding stud. . . . . . . . . . . . . . . . . . 14

16 Cables and connectors side view. The detector signal cables are shown with connectors. Also shown on the lower right is the connector for the thermistor. In the center, on the E-box side is shown the thermostats for the survival heater, and the connector for the thermal hardware. At the lower left are the SEP connector and the grounding stud. . . . . . . . . . . . . . 14

List of Tables

1 Sine Sweep Levels for Instruments Weighing Less than or Equal to 45 kg. Rate = 4 octaves/min. Above 50 Hz, component response may be limited to peak input level. . . . . . . . . . . . 16

2 Random Vibration Levels for Instruments Mounted Directly to Spacecraft Panel and Weighing Less than or Equal to 22.7 kg 16

1 Description of the Housing

The SEPT consists of two identical modules which are mounted at different locations on the spacecraft. One module detects particles in the direction of the solar magnetic field, the other perpendicular to the ecliptic.

Each module is composed of two parts: the housing of the sensor on the top, and the electronics box (E-box) on the bottom. The sensor housing is mounted on top of the E-box with four screws (M4). The sensor is electrically and thermally connected to the E-box.

The E-box is mounted on the spacecraft with four screws (some imperial size, similar to M4). The E-box mounting provides electrical and thermal insulation from the spacecraft. One of the SEPT modules will be mounted on a bracket which is not covered by this note.

Fig. 1 shows the instrument from the front and from the side.

1.1 Materials

The housing is made of aluminum (AA 7075, DIN 3.4365), with a few exceptions:

• The sensor doors are made from titanium, for its low thermal conductance.

• The pinpuller rod is made of bronze, to avoid friction in the door mechanism.

• All screws, washers, etc., are stainless steel.

The material thickness is 1 mm except where noted otherwise.

1.2 Electronics box

The E-box is a rectangular box with removable top and bottom covers.

The body of the box is, like all other parts of the housing, machined from a solid block of aluminum (no welding).

The top (bottom) covers are are fixed with ten (twelve) stainless steel screws (M2) from the sides through holes in the body, into threads with are cut into extension on the covers.

Two Printed Circuit Boards (PCB) with an EMI shield in between are mounted inside the E-box. The PCBs and the EMI shield are assembled outside the box, inserted from the bottom, and screwed with six screws (M3) onto threaded holes that are part of the E-box body. The connectors for the Figure 1: SEPT Housing overview.

5 Figure 2: The body of the electronics box. Figure 3: The electronics box with covers.

Figure 4: The electronics box with bottom (left) and top (right) covers. electronics are pushed from the inside through their respective holes in the E-box body, and fixed with screws.

Figure 5: Sensor top view. Silicon detectors (cyan), magnets (magenta), foil (yellow), foil carrier ring (brown), detector housing (green), Delrin (gray).

1.3 Sensor housing

The sensor (Fig. 5) has four apertures for incident particles, looking in two opposite directions. Each pair of opposite views share a pair of silicon detectors. The particles either have to pass through a thin Parylene foil or through a magnetic field. The foil stops slow protons, and the magnetic field filters electrons.

1.3.1 Magnet assembly

The sensor includes four NdFeB Magnets, size 16×16×8 mm^{3}. The magnets are glued to two yokes, made of Vacoflux 50. The magnet-yoke assemblies are held apart by a magnet support, which also serves to enclose the particle’s path to the detectors. The complete magnet assembly is inserted into the sensor frame, which is 3 mm thick, surrounding the magnet assembly, and which extends with 0.8 mm thick sheets at both sides. See Fig. 6.

The mass of the magnets is 15.6 g each (density 7.63 g/cm^{3}), and each yoke’s mass is 31.2 g (density 8.12 g/cm^{3}).

Figure 6: Magnet assembly. Magnets and yokes (magenta), frame and magnet support (dark blue).

Figure 7: Sensor side support. The side supports slide from below between the extensions of the sensor frame, and the inner support wings of the side supports slide into the slits of the frame and its covers.

1.3.2 Sensor Side support

The magnet assembly will be covered on both sides by frame covers, with round holes for the particle paths. Two side supports will slide between the extensions of the sensor frame. See Figs. 7 and 8. The side support”s inner enforcement wings slide into slits of the sensor frame and cover. The outer support wings also serve to mount pinpullers and a connector.

1.3.3 Collimators

The collimators of the telescope will be mounted on top of the frame covers. Two of these collimator structures also serve as the housing for the detectors and foil assemblies. The hinges for the sensor doors and their release

Figure 8: Sensor side supports and collimator top view.

Figure 9: Collimator assembly. The positions of the screws are indicated on the front view. The grayed screw positions are for the reverse side. Some positions are shared.

mechanism are part of the collimator structure. See Fig. 9.

The sensor is held together by stainless steel screws (M2) which go through the collimators and the frame cover. Four screws from each side go into threads which are cut into the frame, and five go into the magnet support. Four screws go all the way through the side supports and are held by nuts on the other side. Three outermost screws on each side are held with nuts inside the side supports.

1.3.4 Detector assembly

The silicon detector pairs are assembled inside separate housings. The foils are mounted in aluminum rings. A foil carrier rings and a detector housing is mounted inside a collimator. The detector housing is surrounded by Delrin rings, for electrical isolation. See Fig. 5.

1.3.5 The sensor doors

The four sensor openings are covered with doors, which are to be opened only after the spacecraft reached a stable orbit, to prevent direct sunlight to shine on the sensitive detectors and magnets during maneuvers. These doors are not reclosable. The opening mechanism is activated with two pinpullers, opening two doors each. The doors are not sealed.

Fig. 10 shows a side view of the doors in closed an open position, and Fig. 11 shows the front view. The pinpuller, when activated, pulls a rod by a few millimeters and releases the door locks. The doors open by the force of springs.

The left door opens when the rod slips out of its lock, but the right door opens when an extension of the rod slips out of an extended hole in, so that the rod can slip though an opening in the lock. The rod is guided by holes that are part of the collimator structures. These guides are designed such that the (extension of) the rod does not slip out even when the pinpuller is fully activated.

The doors are made of titanium, while the rod is made of bronze. The hinges and guides are part of the collimators which are made of aluminum.

1.4 Connectors and Miscellaneous Attachments

Various objects are attached to the housing.

1.4.1 Pinpullers

The sensor doors are activated by two pinpullers from TiNi Aerospace, Inc., Model P5-403-10S, which are attached with two screws to outer wings of the side supports. These pinpullers weigh 15 g each.

The axis of the pinpuller comes with an internal thread (M2) where the bronze rod is screwed into. The rod diameter is 2 mm. The extension of the shaft that locks the right door is 8 mm long and the extended diameter is 4 mm. See Figs. 12 and 13.

The pinpullers are activated electrically from the spacecraft. The connection is made via a 9-pin subminiature D-type connector which is screwed to an outer wing of a side support. Such a connector weighs 6 g.

Figure 10: Sensor doors side view. Left: doors in closed position, right: doors in open position.

Figure 11: Sensor doors front view, with the doors opened. The pinpuller rod is drawn in both the idle (dark blue) and activated (pink) position.

Figure 12: Pinpuller mounting top view. Also shown is the connector for the supply of the pinpuller activation current.

Figure 13: Pinpuller mounting side view. Only the near half of the sensor is shown. The sensor doors are drawn both opened and closed.

1.4.2 Purge nipple

To purge the sensor with nitrogen gas before launch there is a purge nipple screwed into the sensor frame from the side. The side support has hole at that position. The design of the purge nipple is not yet final, because a flow limiter needs to be incorporated. See Fig. 14.

1.4.3 Detector signals

The detector signals are routed through twelve coax cables type RG 178, six for each detector pair, which leave the detector housing to the side, over an angle of 135^{∘}. The cables attach to the E-box via SSMC type connectors.

Figure 14: Purge nipple and thermistor (green). The purge nipple is the upper left and the thermistor is the lower right part in green.

Such a connector pair weighs 2.5 g.

The cable will need to be fixed to the E-box cover somehow. This part of the design is not yet finished.

Figs. 15 and 16 show the positions of these cables and various other connectors.

1.4.4 SEP connection

The electronics upstream connection of the E-box to SEP central goes through a 31-pin MDM connector.

1.4.5 Thermistor

A thermistor is required to measure the temperature of the sensor, close to the silicon detectors. The thermistor is glued into an stainless steel cylinder, which is attached to a side support from the inner side below the sensor. The connection is made with a coax cable from the inner end of the thermistor to a connector on the back side of E-box, opposite to the SEP connector. Cable and connector are the same type as for the detector signals. See Fig. 14.

1.4.6 Thermal hardware

Survival and operational heaters will be attached on the outside to the bottom of the E-box. Two thermostats for the survival heater will be glued to the side of the E-box. The electrical connections to the spacecraft will use a 9-pin MDM connector which is attached to the E-box near the thermostats. See Figs. 15 and 16.

Figure 15: Cables and connectors front view. The detector signal cables are drawn in light green. The connector on the right between the support wings of the E-box supplies the thermal hardware. The MDM connector on the front is the link to SEP central. To its left is the grounding stud.

Figure 16: Cables and connectors side view. The detector signal cables are shown with connectors. Also shown on the lower right is the connector for the thermistor. In the center, on the E-box side is shown the thermostats for the survival heater, and the connector for the thermal hardware. At the lower left are the SEP connector and the grounding stud.

1.4.7 Ground Stud

A stainless steel grounding stud is screwed to the front of the E-box next to the SEP connector. See Figs. 15 and 16.

2 Vibration levels

These are quoted from the environmental specifications for instruments on the STEREO spacecraft.

2.1 Instrument Stiffness

Instruments shall be designed such that the primary vibration modes are above 50 Hz.

2.2 Instrument Sine Sweep Vibration Tests

Instruments shall be subjected to the following sinusoidal vibration levels. These shall be applied in each of three orthogonal axes.

See Table 1 for the levels.

2.3 Random Vibration Levels

All instruments shall be subjected to the following random vibration levels. These shall be applied to each of three orthogonal axes, one of which is parallel to the thrust axis.

See Table 2 for the levels.

Table 1: Sine Sweep Levels for Instruments Weighing Less than or Equal to 45 kg. Rate = 4 octaves/min. Above 50 Hz, component response may be limited to peak input level.

Thrust Axes (X)

Frequency Acceleration
5 to 7.4 Hz  0.5 in (double amplitude)
7.4 to 23 Hz  1.4 g
25 to 27 Hz  16.0 g
29 to 100 Hz  1.4 g

Lateral Axes (Y,Z)

Frequency Acceleration
5 to 6.3 Hz  0.5 in (double amplitude)
6.3 to 19 Hz  1.0 g
21 to 23 Hz  12.0 g
25 to 100 Hz  1.0 g

Table 2: Random Vibration Levels for Instruments Mounted Directly to Spacecraft Panel and Weighing Less than or Equal to 22.7 kg

Perpendicular to Mounting Panel

Frequency  PSD Level
20 Hz  0.0063 g^{2}/Hz
20 to 80 Hz  +6 dB/oct
80 to 800 Hz  0.1 g^{2}/Hz
800 to 2000 Hz  −9 dB/oct
2000 Hz  0.0065 g^{2}/Hz

Overall Amplitude = 10.4 g rms Duration = 60 seconds

Parallel to Mounting Panel

Frequency  PSD Level
20 Hz  0.0031 g^{2}/Hz
20 to 80 Hz  +6 dB/oct
80 to 800 Hz  0.05 g^{2}/Hz
800 to 2000 Hz  −9 dB/oct
2000 Hz  0.0032 g^{2}/Hz

Overall Amplitude = 7.4 g rms